Photoelectrons with a maximum speed of 7.00 · 105 m/sec are ejected from a surface in the presence of light with a frequency of 8.00 · 1014Hz. Calculate the energy required to eject photoelectrons from the surface. _______a0 x 10a1 _______joules
Answer #1

Energy required,E=(hv) - ( .5 mv^2 )= (528*10^(-21)) - (.5 * (9.1*10^(-31))*49*10^10) = 305.05 * 10^(-21) joule